# Differential Machine Learning for 0DTE Options with Stochastic Volatility and Jumps

### 08 Mar 2026 v1

@@ Abstract
We present a differential machine learning method for zero-days-to-expiry (0DTE) options under a stochastic-volatility jump-diffusion model that computes prices and Greeks in a single network evaluation. To handle the ultra-short-maturity regime, we represent the price in Black–Scholes form with a maturity-gated variance correction, and combine supervision on prices and Greeks with a PIDE-residual penalty. To make the jump contribution identifiable, we introduce a separate jump-operator network and train it with a three-stage procedure. In Bates-model simulations, the method improves jump-term approximation relative to one-stage baselines, keeps price errors close to one-stage alternatives while improving Greeks accuracy, produces stable one-day delta hedges, and is substantially faster than a Fourier-based pricing benchmark.
@@ 1 Introduction
Zero-days-to-expiry (0DTE) options have grown rapidly over the past few years and account for a large fraction of trading volume. Recent empirical work finds elevated intraday jump activity and large gamma exposures around the at-the-money (ATM) region [Bandi et al.(2023), Bozovic(2025), Dim et al.(2024)]. These features raise two practical challenges: (i) the underlying dynamics are better captured by a diffusion model with jumps, and (ii) the very short maturities and frequent intraday rebalancing require fast computation of option prices and Greeks.
We apply differential machine learning (DML)[Huge and Savine(2020)] to 0DTE options under the stochastic-volatility jump-diffusion (SVJD) model. Learning-based models can reduce pricing cost: once trained, prices and Greeks are obtained from a single network evaluation. However, the numerically most unstable region is near the money, where Greeks can become very large for 0DTE options with jumps.
A common approach in machine-learning PDE solvers is to enforce the governing PDE by penalizing its residual at sampled state points, together with terminal/boundary conditions [Raissi et al.(2019)]. Several recent studies apply this residual-penalty approach to option pricing with jumps by including a PIDE-residual term in the training objective [Fu and Hirsa(2020), Sun et al.(2025), Bansal et al.(2026)]. Three design choices in our scheme are central:
@@
    We adopt DML which trains a single price network on both option values and Greeks. Greeks are computed by automatic differentiation of the network output with respect to its inputs and are included directly in the training loss function.
@@
    Instead of predicting prices directly, the network learns a variance correction in a Black–Scholes formula[Black and Scholes(1973)], scaled so that the correction vanishes as $\tau\to 0$. This ensures the correct short-maturity limit (payoff) and reduces the approximation difficulty faced by the network in the near-singular regime. The approach is common in learning-based volatility surface models [Liu et al.(2019), Ackerer et al.(2020)].
@@
    We introduce a second neural network to represent the compensated jump operator. When the jump component is identified only through a PIDE-residual penalty, the optimizer can trade off errors between the diffusion and jump terms while keeping the overall residual small. As a result, a small residual does not imply that the learned jump operator matches the model-implied jump integral. Therefore we use the second network to make the residual penalty informative about the jump term.
@@
Section 2 presents the Bates model. Section 3 introduces the DML-based neural network and the three-stage training scheme. Section 4 specifies the loss functions and constraints. Section 5 reports numerical experiments, and the appendix provides additional results. All quantitative results in this paper are based on simulated data generated from the specified stochastic models.
@@ 2 Bates model
We work under the Bates stochastic-volatility jump-diffusion model, which combines a Heston-type variance process with Merton-style lognormal price jumps [Merton(1976), Heston(1993), Bates(1996)]. Under the risk-neutral measure let $S_{t}$ denote the underlying asset, $V_{t}$ the instantaneous variance, $r$ the risk-free rate, and $q$ the dividend. The dynamics are
@@
For a European call option with maturity $\tau$ and strike $K$, the risk-neutral call price $C$ solves the following PIDE
@@ 3 DML for PIDEs
Figure 1 summarizes our neural-network model. The solid arrows correspond to forward evaluations while dashed arrows indicate computations derived from these outputs. The input is
@@
and return a Black–Scholes call price with volatility $\sigma_{\mathrm{eff}}=\sqrt{V_{\mathrm{eff}}}$. The $g(\tau)$ forces the learned variance correction to vanish near expiry, which stabilizes training and preserves the payoff limit. Since we work in log-moneyness, we normalize the strike to $K=1$:
@@
Figure 1: Architecture and training procedure. Stages 1–3 describe the three-stage training scheme. A variance-correction network returns $\Delta V_{\phi}(\mathbf{x})$, multiplied by a deterministic maturity function $g(\tau)$ so that the correction vanishes as $\tau\to 0$. Prices are produced by substituting the resulting effective variance into the Black–Scholes call formula, yielding $u_{\phi}(\mathbf{x})$ (not the plain BS price unless $\Delta V_{\phi}\equiv 0$). A separate network outputs the compensated jump contribution $J_{\psi}(\mathbf{x})$. Greeks are obtained by automatic differentiation of $u_{\phi}$. The jump-PIDE residual $R(\mathbf{x})$ is computed and penalized at randomly sampled points. Stages 1–3 depict the three-stage training scheme used for jump-term identifiability.
3.1 Twin-network
A standard model fits a network $u_{\phi}(\mathbf{x})$ by
@@ 3.1 Twin-network
Training can be further regularized by adding a PDE-residual penalty computed from the same automatic-differentiation derivatives. But if we approximate the nonlocal jump operator $J[u]$ by a separate network $J_{\psi}(\mathbf{x})$ and use it only through the residual
@@
$u_{\phi}$ and $J_{\psi}$ are not separately identifiable: in principle, for any $u_{\phi}$ one can set $J_{\psi}=-\mathcal{L}_{\mathrm{diff}}u_{\phi}/\lambda$ and obtain $R\equiv 0$. This mechanism can lead to a degenerate solution in which the jump network cancels diffusion-operator errors and reduces the residual without learning a meaningful jump contribution. We therefore supervise the jump operator explicitly. Using the reference price $u^{\text{ref}}$, we construct a numerical proxy for the jump term and train $J_{\psi}$ to match it.
@@ 3.2 Three-stage training with jump supervision
    Stage 1 (price and Greeks). Train the price network (freeze $J_{\psi}$) using the price and Greek terms plus the no-arbitrage penalties.
@@
    On randomly sampled state points we approximate the expectation in (12) with a Gauss–Hermite rule applied to $u^{\text{ref}}(x+Y,V,\tau)$ and penalize $J_{\psi}-J^{\text{ref}}(x,V,\tau)$ using a Huber loss function.
@@ 4.1 Loss function, constraints, and weights
where $(z_{i},w_{i})_{i=1}^{n_{\mathrm{SC}}}$ are nodes/weights of a Gauss–Hermite rule for $Z\sim\mathcal{N}(0,1)$ and we use $\omega_{\mathrm{SC}}=0.05,n_{\mathrm{SC}}=16$.
@@ 5 Numerical experiments
Our experiments address two questions: How much do DML and residual regularization improve 0DTE price/Greek accuracy, and do the resulting prices and Greeks remain reliable in simple hedging tests? Additional model-choice robustness checks (BS/Merton baselines and an SVCJ extension) are in Appendix A.
@@ 5.1 Setting
  • Log-moneyness:
    with probability $p_{\text{core}}=0.7$ we draw from a near-ATM band $x\sim\mathrm{Unif}[-0.1,0.1]$ (“ATM core”);
    otherwise $x\sim\mathrm{Unif}[-0.5,0.5]$.
  • Maturity (strict 0DTE):
    with probability $p_{\text{short}}=0.7$ we draw very short maturities $\tau\sim\mathrm{Unif}[10^{-4},1/504]$;
    otherwise $\tau\sim\mathrm{Unif}[1/504,1/252]$. Here $p_{\text{short}}$ is the mixture weight used to oversample the shortest maturities.
@@
To stabilize jump-term training, we add additional state points by pushing data points through the jump map $x\mapsto x+Y$ and clamping to $|x|\leq 6$. We compute reference prices on these jump-shifted (“margin”) points and add a small auxiliary price-consistency loss. This extends supervision beyond $|x|\leq 0.5$ to the region visited by the jump integral. Deep OTM options have tiny prices and noisy Greeks, so we ignore Greek targets when the reference price is below $10^{-4}$.
We use $\omega_{G}=0.7$, $\omega_{R}=0.01$, $\omega_{\text{col}}=0.1$, $\omega_{\text{NA}}=0.05$, and $(\omega_{\Delta},\omega_{\Gamma},\omega_{\nu})=(1.0,0.3,0.5)$. We also set the self-consistency weight $\omega_{\mathrm{SC}}=0.05$ with $n_{\mathrm{SC}}=16$. We set $(W_{\text{ATM}},W_{\text{SHORT}},W_{\text{ATM\&SHORT}})=(1.0,1.0,1.0)$.
@@ 5.2 Accuracy across models
We train the one-stage variants (Models A–C) for 100 epochs with Adam (learning rate $10^{-3}$) and batch size 64. The three-stage model uses 100 epochs (Stage 1), 60 epochs (Stage 2), and 30 epochs (Stage 3). Table 1 reports RMSE and Table 2 reports RMSE in a region $|x|<0.05$ and $\tau\leq 1/252$. Residual statistics are in Table 3.
@@
Table 4: Normalised RMSEs and tail absolute errors (validation sample; $N=500$, seed=42).
@@
We additionally report scale-invariant metrics and tail error quantiles in Table 4 with two scale-free versions of the price RMSE:
@@
The tail metrics $|e_{P}|_{p}$ and $|e_{\Gamma}|_{p}$ denote the $p$-th percentile of the absolute price and gamma errors, respectively. Tables 14 suggest three points. First, Model B improves delta accuracy (first-order risk) and vega accuracy relative to the price-only fit (Model A) (Table 1). For gamma $\Gamma$, improvements are clearer in tail error quantiles than in global RMSE (Table 4). Second, the one-stage PIDE residual penalty (Model C) does not improve prices or Greeks relative to Model B. This is consistent with jump settings, where the residual can be reduced via cancellation between the differential and jump terms rather than by learning an interpretable jump operator. Third, the three-stage model (Model D) yields the best overall $\Delta$ and vega RMSE among the Greek-supervised models and slightly improves $\Gamma$ tail error quantiles, although $\Gamma$ remains the most delicate target in the region $(|x|<0.05,\,\tau\leq 1/252)$.
Figure 2 visualizes true vs predicted prices and Greeks for Model D. In the ultra-short maturity regime ($\tau\to 0$), $\Gamma$ is extremely localised around the money: for the majority of (deep ITM/OTM) evaluation points the true $\Gamma$ is numerically close to zero, so the scatter plots naturally show a dense cluster at $\Gamma\approx 0$.
@@ 5.3 Jump-term comparison
This subsection clarifies why a small PIDE residual is not evidence that the learned jump component represents the intended compensated jump integral. We first report error distributions for the final three-stage model (Figure 3), and then directly test whether the learned compensated jump contribution $J_{\psi}$ matches a numerical proxy computed from the reference pricer (Figure 4).
@@ 5.5 One-day delta-hedging
The data-generating process is the Bates SVJD model and we simulate $n_{\text{paths}}=5000$ paths of $(S_{t},V_{t})$ over one trading day $T=1/252$ using an Euler scheme with $n_{\text{steps}}=24$. Jumps are simulated by a Bernoulli approximation with step probability $p\approx\lambda\,\Delta t$. To avoid unrealistically large per-step jump probabilities when $\lambda\,\Delta t$ is not very small, we cap this probability at $0.2$, i.e. we use $p=\min(\lambda\,\Delta t,0.2)$. We use a stressed parameter set
@@
chosen to amplify jump activity and gamma effects. First we conduct stock-only $\Delta$ hedges: the hedge portfolio contains the underlying and a cash account. We run the hedge separately for each strike $K\in\{0.9,1.0,1.1\}$ on the same set of simulated $(S_{t},V_{t})$ paths, which produces 5000 P&L outcomes per strike. At $t=0$ we short one call with strike $K\in\{0.9,1.0,1.1\}$ and maturity $T$. At each time $t_{k}$ we compute delta from either the Fourier pricer (true) or the model (DML), rebalance the stock position, and carry the residual in a cash account accruing at rate $r$. At maturity we record hedging P&L as final portfolio value minus option payoff. Aggregating across strikes ($15000$ paths total), we report the summary statistics in Table 6. Figure 5 plots P&L histograms which show that the distributional difference between true and model delta hedges is small.
@@
Second we compare three one-option second-order hedges: (i) a ratio-type $\Delta+\Gamma$ hedge (ii) a weighted ridge least-squares (LS) hedge that fits $(\Delta,\Gamma)$ jointly across stock and one hedge option, and (iii) a P&L-increment regression hedge (PL-LS) that learns state-dependent hedge ratios with ridge regularization.
@@
and at each rebalance date update $(w_{\mathrm{H},t},w_{S,t})$ and the cash account so that the strategy remains self-financing. The weighted ridge LS hedge uses the same instrument pair $(S_{t},C^{\mathrm{H}}_{t})$, but chooses $(w_{S},w_{\mathrm{H}})$ by minimising
@@
PL-LS estimates time-dependent coefficient vectors $(A_{k},B_{k})$ from a ridge regression of option price increments $\Delta C_{\text{main}}$ on basis-weighted $(\Delta S,\Delta C_{\text{hedge}})$, and then sets the hedge with $w_{S}=\phi^{\top}A_{k}$ and $w_{\mathrm{H}}=\phi^{\top}B_{k}$ (with $\phi=[1,\log m,(\log m)^{2},V,\tau]^{\top}$).
We consider an option fixed at a one-day ATM call with $K_{\mathrm{main}}=1.00$ and $T=1/252$. We then run four separate hedging experiments, each using the stock plus one additional same-maturity call with hedge strike $K_{\mathrm{H}}\in\{0.99,1.00,1.01,1.02\}$. Figure 6 pools the resulting P&L observations across these four separate $K_{\mathrm{H}}$ choices.
@@
The right panel shows the tail of the ratio-type $\Delta+\Gamma$ hedge via the CCDF of $|\mathrm{P\&L}|$. LS/PL-LS are omitted on the right because their tails collapse near zero at this scale.
Figure 6: $\Delta+\Gamma$ hedging illustration for a one-day ATM liability call ($K_{\mathrm{main}}=1.00$). In each run the hedge portfolio contains the stock and one additional one-day call at a single strike $K_{\mathrm{H}}\in\{0.99,1.00,1.01,1.02\}$; the four $K_{\mathrm{H}}$ choices are run separately and the plotted distributions pool the resulting outcomes across them. Left: central-region density of hedging P&L for ratio-type $\Delta+\Gamma$, LS, and PL-LS, shown separately for (top) true Greeks and (bottom) DML Greeks. The two central-density panels are almost indistinguishable at this scale, reflecting that the DML Greeks are close to the benchmark. Right: tail behaviour of the ratio-type $\Delta+\Gamma$ hedge, shown as the CCDF of $|\mathrm{P\&L}|$ (true vs. DML) on log scales; a one-option gamma-based ratio hedge can become unstable when the hedge option’s gamma is near zero.

### 30 Mar 2026 v2

@@ Abstract
We present a differential machine learning method for zero-days-to-expiry (0DTE) options under a stochastic-volatility jump-diffusion model that computes prices and Greeks in a single network evaluation. To handle the ultra-short-maturity regime, we represent the price in Black–Scholes form with a maturity-gated variance correction and combine supervision on prices and Greeks with a PIDE-residual penalty. To make the jump contribution identifiable, we introduce a separate jump-operator network and train it with a three-stage procedure. In Bates-model simulations, the method improves jump-term approximation relative to one-stage baselines, keeps price errors comparable, improves Greek accuracy, produces stable one-day delta hedges, and is substantially faster than a Fourier-based pricing benchmark.
@@ 1 Introduction
Zero-days-to-expiry (0DTE) options have grown rapidly over the past few years and account for a large fraction of trading volume. Recent empirical work documents frequent intraday jumps in 0DTE markets [Bozovic(2025)], studies the role of gamma-related open-interest exposures [Dim et al.(2024)], and analyzes the pricing implications of short-horizon tail risk in 0DTE options [Bandi et al.(2023)]. These findings raise two practical challenges: (i) the underlying dynamics are better captured by a diffusion model with jumps, and (ii) the very short maturities and frequent intraday rebalancing require fast computation of option prices and Greeks.
We apply differential machine learning (DML) [Huge and Savine(2020), Frandsen et al.(2022)] to 0DTE options under the stochastic-volatility jump-diffusion (SVJD) model. Learning-based models can reduce pricing costs: once trained, prices and Greeks are obtained from a single network evaluation. However, the numerically most challenging region is near the money, where Greeks can become very large for 0DTE options with jumps.
A common approach in machine-learning PDE solvers is to enforce the governing PDE by penalizing its residual at sampled state points, together with terminal and boundary conditions [Raissi et al.(2019)]. Several recent studies apply this residual-penalty approach to option pricing with jumps by including a PIDE-residual term in the training objective [Fu and Hirsa(2020), Sun et al.(2025), Bansal et al.(2026)]. Three design choices are central to our approach:
@@
    We adopt DML, which trains a single price network on both option values and Greeks. Greeks are computed by automatic differentiation of the network output with respect to its inputs and are included directly in the training loss function.
@@
    Instead of predicting prices directly, the network learns a variance correction in a Black–Scholes formula [Black and Scholes(1973)], scaled so that the correction vanishes as $\tau\to 0$. This ensures the correct short-maturity limit (payoff) and reduces the approximation difficulty faced by the network in the near-singular regime. More broadly, this design is in the spirit of learning-based option-pricing and implied-volatility models that impose financial structure to improve stability and surface regularity [Liu et al.(2019), Ackerer et al.(2020)].
@@
    We introduce a second neural network to represent the compensated jump operator. When the jump component is identified only through a PIDE-residual penalty, the optimizer can trade off errors between the diffusion and jump terms while keeping the overall residual small. As a result, a small residual does not imply that the learned jump operator matches the model-implied jump integral. Therefore, we use the second network to make the residual penalty informative about the jump term.
@@
Section 2 presents the Bates model. Section 3 introduces the DML-based neural network and the three-stage training scheme. Section 4 specifies the loss functions and constraints. Section 5 reports numerical experiments, and the appendix provides additional results. All quantitative results are based on simulated data generated from the specified stochastic models.
@@ 2 Bates model
We work with the Bates stochastic-volatility jump-diffusion model, which combines a Heston-type variance process with Merton-style lognormal price jumps [Merton(1976), Heston(1993), Bates(1996)]. Under the risk-neutral measure, let $S_{t}$ denote the underlying asset, $V_{t}$ the instantaneous variance, $r$ the risk-free rate, and $q$ the dividend. The dynamics are
@@
For a European call option with maturity $\tau$ and strike $K$, the risk-neutral call price $C$ satisfies the following PIDE:
@@ 3 DML for PIDEs
Figure 1 summarizes the model architecture. The solid arrows correspond to forward evaluations while dashed arrows indicate computations derived from these outputs. The input is
@@
and we return a Black–Scholes call price with volatility $\sigma_{\mathrm{eff}}=\sqrt{V_{\mathrm{eff}}}$. The $g(\tau)$ forces the learned variance correction to vanish near expiry, which stabilizes training and preserves the payoff limit. Since we work in log-moneyness, we normalize the strike to $K=1$:
@@
Figure 1: Architecture and training procedure. Stages 1–3 describe the three-stage training scheme. A variance-correction network returns $\Delta V_{\phi}(\mathbf{x})$, multiplied by a deterministic maturity function $g(\tau)$ so that the correction vanishes as $\tau\to 0$. Prices are produced by substituting the resulting effective variance into the Black–Scholes call formula, yielding $u_{\phi}(\mathbf{x})$ (not the standard BS price unless $\Delta V_{\phi}\equiv 0$). A separate network outputs the compensated jump contribution $J_{\psi}(\mathbf{x})$. Greeks are obtained by automatic differentiation of $u_{\phi}$. The jump-PIDE residual $R(\mathbf{x})$ is computed and penalized at randomly sampled points. Stages 1–3 depict the three-stage training scheme used for jump-term identifiability.
3.1 Twin network
A standard supervised model fits a network $u_{\phi}(\mathbf{x})$ by
@@ 3.1 Twin network
Training can be further regularized by adding a PDE-residual penalty computed from the same automatic-differentiation derivatives. However, if we approximate the nonlocal jump operator $J[u]$ by a separate network $J_{\psi}(\mathbf{x})$ and use it only through the residual
@@
$u_{\phi}$ and $J_{\psi}$ are not separately identifiable: in principle, for any $u_{\phi}$ one can set $J_{\psi}=-\mathcal{L}_{\mathrm{diff}}u_{\phi}/\lambda$ and obtain $R\equiv 0$. This mechanism can lead to a degenerate solution in which the jump network cancels diffusion-operator errors and reduces the residual without learning a meaningful jump contribution. We therefore supervise the jump operator explicitly. Using reference prices $u^{\text{ref}}$, we construct a numerical proxy for the jump term and train $J_{\psi}$ to match it.
@@ 3.2 Three-stage training with jump supervision
    Stage 1 (price and Greeks). Train the price network while freezing $J_{\psi}$, using the price and Greek terms plus the no-arbitrage penalties.
@@
    On randomly sampled state points, we approximate the expectation in (12) with a Gauss–Hermite rule applied to $u^{\text{ref}}(x+Y,V,\tau)$ and penalize $J_{\psi}-J^{\text{ref}}(x,V,\tau)$ using a Huber loss function.
@@ 4.1 Loss function, constraints, and weights
where $(z_{i},w_{i})_{i=1}^{n_{\mathrm{SC}}}$ are nodes and weights of a Gauss–Hermite rule for $Z\sim\mathcal{N}(0,1)$, and we use $\omega_{\mathrm{SC}}=0.05,n_{\mathrm{SC}}=16$.
@@ 5 Numerical experiments
Our experiments address two questions: how much DML and residual regularization improve 0DTE price/Greek accuracy, and whether the resulting prices and Greeks remain reliable in simple hedging tests. Additional model-choice robustness checks (BS/Merton baselines and an SVCJ extension) are reported in Appendix A.
@@ 5.1 Setting
  • Log-moneyness: with probability $p_{\text{core}}=0.7$ we draw from a near-ATM band $x\sim\mathrm{Unif}[-0.1,0.1]$ (“ATM core”); otherwise $x\sim\mathrm{Unif}[-0.5,0.5]$.
  • Maturity (strict 0DTE): with probability $p_{\text{short}}=0.7$ we draw very short maturities $\tau\sim\mathrm{Unif}[10^{-4},1/504]$; otherwise $\tau\sim\mathrm{Unif}[1/504,1/252]$. Here $p_{\text{short}}$ is the mixture weight used to oversample the shortest maturities.
@@
To stabilize jump-term training, we augment the sampled states by pushing points through the jump map $x\mapsto x+Y$ and clamping to $|x|\leq 6$. We compute reference prices on these jump-shifted (“margin”) points and add a small auxiliary price-consistency loss. This extends supervision beyond $|x|\leq 0.5$ to the region visited by the jump integral. Deep OTM options have tiny prices and noisy Greeks, so we ignore Greek targets when the reference price is below $10^{-4}$.
We use $\omega_{G}=0.7$, $\omega_{R}=0.01$, $\omega_{\text{col}}=0.1$, $\omega_{\text{NA}}=0.05$, and $(\omega_{\Delta},\omega_{\Gamma},\omega_{\nu})=(1.0,0.3,0.5)$. We also set the self-consistency weight $\omega_{\mathrm{SC}}=0.05$ with $n_{\mathrm{SC}}=16$. Finally, we set $(W_{\text{ATM}},W_{\text{SHORT}},W_{\text{ATM\&SHORT}})=(1.0,1.0,1.0)$.
@@ 5.2 Accuracy across models
We train the one-stage variants (Models A–C) for 100 epochs with Adam (learning rate $10^{-3}$) and batch size 64. The three-stage model uses 100 epochs (Stage 1), 60 epochs (Stage 2), and 30 epochs (Stage 3). Table 1 reports global RMSE, and Table 2 reports RMSE in the region $|x|<0.05$ and $\tau\leq 1/252$. Residual statistics are in Table 3.
@@
Table 4: Normalized RMSEs and tail absolute errors (validation sample; $N=500$, seed=42).
@@
We additionally report scale-invariant metrics and tail error quantiles in Table 4. For price RMSE, we use two scale-free metrics:
@@
The tail metrics $|e_{P}|_{p}$ and $|e_{\Gamma}|_{p}$ denote the $p$-th percentile of the absolute price and gamma errors, respectively. Tables 14 suggest three points. First, Model B improves delta accuracy (first-order risk) and vega accuracy relative to the price-only fit (Model A) (Table 1). For gamma $\Gamma$, improvements are clearer in tail error quantiles than in global RMSE (Table 4). Second, the one-stage PIDE residual penalty (Model C) does not improve prices or Greeks relative to Model B. This is consistent with jump models, where the residual can be reduced via cancellation between the differential and jump terms rather than by learning an interpretable jump operator. Third, the three-stage model (Model D) yields the best overall $\Delta$ and vega RMSE among the Greek-supervised models and slightly improves $\Gamma$ tail error quantiles, although $\Gamma$ remains the most delicate target in the region $(|x|<0.05,\,\tau\leq 1/252)$.
Figure 2 visualizes true vs predicted prices and Greeks for Model D. In the ultra-short maturity regime ($\tau\to 0$), $\Gamma$ is extremely localized around the money: for the majority of (deep ITM/OTM) evaluation points the true $\Gamma$ is numerically close to zero, so the scatter plots naturally show a dense cluster at $\Gamma\approx 0$.
@@ 5.3 Jump-term comparison
This subsection clarifies why a small PIDE residual alone is not evidence that the learned jump component represents the intended compensated jump integral. We first report error distributions for the final three-stage model (Figure 3), and then directly test whether the learned compensated jump contribution $J_{\psi}$ matches a numerical proxy computed from the reference pricer (Figure 4).
@@ 5.5 One-day delta-hedging
The data-generating process follows the Bates SVJD model, and we simulate $n_{\text{paths}}=5000$ paths of $(S_{t},V_{t})$ over one trading day $T=1/252$ using an Euler scheme with $n_{\text{steps}}=24$. Jumps are simulated by a Bernoulli approximation with step probability $p\approx\lambda\,\Delta t$. To avoid unrealistically large per-step jump probabilities when $\lambda\,\Delta t$ is not very small, we cap this probability at $0.2$, i.e. we use $p=\min(\lambda\,\Delta t,0.2)$. We use a stressed parameter set
@@
chosen to amplify jump activity and gamma effects. First, we conduct stock-only $\Delta$ hedges, where the hedge portfolio contains the underlying and a cash account. We run the hedge separately for each strike $K\in\{0.9,1.0,1.1\}$ on the same set of simulated $(S_{t},V_{t})$ paths, which produces 5000 P&L outcomes per strike. At $t=0$, we short one call with strike $K\in\{0.9,1.0,1.1\}$ and maturity $T$. At each time $t_{k}$ we compute delta from either the Fourier pricer (true) or the model (DML), rebalance the stock position, and carry the residual in a cash account accruing at rate $r$. At maturity we record hedging P&L as final portfolio value minus option payoff. Aggregating across strikes ($15000$ paths total), we report summary statistics in Table 6. Figure 5 plots P&L histograms, showing that the distributional difference between true and model delta hedges is small.
@@
Second, we compare three one-option second-order hedges: (i) a ratio-type $\Delta+\Gamma$ hedge; (ii) a weighted ridge least-squares (LS) hedge that fits $(\Delta,\Gamma)$ jointly across stock and one hedge option; and (iii) a P&L-increment regression hedge (PL-LS) that learns state-dependent hedge ratios with ridge regularization.
@@
and at each rebalance date update $(w_{\mathrm{H},t},w_{S,t})$ and the cash account so that the strategy remains self-financing. The weighted ridge LS hedge uses the same instrument pair $(S_{t},C^{\mathrm{H}}_{t})$, but chooses $(w_{S},w_{\mathrm{H}})$ by minimizing
@@
PL-LS estimates time-dependent coefficient vectors $(A_{k},B_{k})$ from a ridge regression of option price increments $\Delta C_{\text{main}}$ on basis-weighted $(\Delta S,\Delta C_{\text{hedge}})$ and then sets the hedge using $w_{S}=\phi^{\top}A_{k}$ and $w_{\mathrm{H}}=\phi^{\top}B_{k}$ (with $\phi=[1,\log m,(\log m)^{2},V,\tau]^{\top}$).
We fix the liability option as a one-day ATM call with $K_{\mathrm{main}}=1.00$ and $T=1/252$. We then run four separate hedging experiments, each using the stock plus one additional same-maturity call with hedge strike $K_{\mathrm{H}}\in\{0.99,1.00,1.01,1.02\}$. Figure 6 pools the resulting P&L observations across these four $K_{\mathrm{H}}$ choices.
@@
The right panel shows the tail of the ratio-type $\Delta+\Gamma$ hedge via the CCDF of $|\mathrm{P\&L}|$. LS and PL-LS are omitted on the right because their tails collapse near zero at this scale.
Figure 6: $\Delta+\Gamma$ hedging illustration for a one-day ATM liability call ($K_{\mathrm{main}}=1.00$). In each run the hedge portfolio contains the stock and one additional one-day call at a single strike $K_{\mathrm{H}}\in\{0.99,1.00,1.01,1.02\}$; the four $K_{\mathrm{H}}$ choices are run separately and the plotted distributions pool the resulting outcomes across them. Left: central-region density of hedging P&L for ratio-type $\Delta+\Gamma$, LS, and PL-LS, shown separately for (top) true Greeks and (bottom) DML Greeks. The two central-density panels are almost indistinguishable at this scale, reflecting that the DML Greeks are close to the benchmark. Right: tail behavior of the ratio-type $\Delta+\Gamma$ hedge, shown as the CCDF of $|\mathrm{P\&L}|$ (true vs. DML) on log scales; a one-option gamma-based ratio hedge can become unstable when the hedge option’s gamma is near zero.